# SFY — A lightweight, high-frequency and phase-resolving wave-buoy for coastal waters

### v2

@@ 6.2 Capturing and quantifying breaking waves
In the open water (Sec. 4) the breaking fraction is quantified using methods developed by Brown et al. (2018). The fraction qualitatively agrees with photos from the location. However, a more controlled experiment with camera on the buoy, or the buoy located in the field of vision of stereo-video cameras, would provide more reliable context. This would allow better evaluation of how robust this method is and the influence of the parameters that go into the it.
@@
The reason that the buoy can measure this way is presumably due to a combination of using a small wave buoy, high frequency sampling, and a wide measure range on the IMU (accelerometer and gyroscope).
@@ 7 Conclusions
The SFY, a small, light-weight, drifting or moored wave buoy design with the capability to detect breaking events has been described. The buoy is designed to work in coastal environments and communicates over the cellular network as well as stores the data on SD-card. Its design, code, and hardware is open source and available publicly [^8] https://github.com/gauteh/sfy. The buoy has been proven to work reliably up to $80$ km from the nearest cellular tower, and in its small form measures frequencies from about $0.04$ Hz to about $2.2$ Hz (or $1.1$ Hz for the moored buoy, without requiring correction). The main advantages and improvements of the SFY buoy are:

### v3

@@ 6.2 Capturing and quantifying breaking waves
In the open water (Sec. 4) the breaking fraction is quantified using methods developed by Brown et al. (2018). The fraction qualitatively agrees with photos from the location. However, a more controlled experiment with camera on the buoy, or the buoy located in the field of vision of stereo-video cameras, would provide more reliable context. This would allow better evaluation of how robust this method is and the influence of the parameters that go into it.
@@
The reason that the buoy can measure this way is presumably due to a combination of using a small wave buoy, high frequency sampling, and a wide measurement range on the IMU (accelerometer and gyroscope).
@@ 7 Conclusions
The SFY, a small, light-weight, drifting or moored wave buoy design with the capability to detect breaking events has been described. The buoy is designed to work in coastal environments and communicates over the cellular network as well as storing the data on SD-card. Its design, code, and hardware is open source and available publicly [^9] https://github.com/gauteh/sfy. The buoy has been proven to work reliably up to $80$ km from the nearest cellular tower, and in its small form measures frequencies from about $0.04$ Hz to about $2.2$ Hz (or $1.1$ Hz for the moored buoy, without requiring correction). The main advantages and improvements of the SFY buoy are: